OpenStudy (anonymous):

y= x ² - 6x + 8 What are the X intercepts and Y intercept

OpenStudy (anonymous):

is that a 2 in the superscript?

OpenStudy (anonymous):

Its X squared

OpenStudy (anonymous):

x intercept,x=2,4

OpenStudy (anonymous):

x^2-6x+8=(x-4)(x-2) To find x intercepts , suppose y=0 , y=(x-4)(x-2) so (x-4)(x-2)=0 , x intercepts are 4 and 2 To find y intercept , plug in 0 to x so -4*-2=8

OpenStudy (anonymous):

y intercept,y=8

OpenStudy (anonymous):

Thanks guys